OpenStudy (anonymous):

how to simplify 64^3/2?

OpenStudy (raden):

hint : 64 = 8^2 then use the property: (a^m)^n = a^(m*n)

OpenStudy (aayushi.somani):

see \[\frac{ (8^{2})^{3} }{ 2 }\] \[\frac{ 8^{6} }{ 2 }\] now u can solve it..

OpenStudy (raden):

@memirobb , you are asking about 64^(3/2) or (64^3)/2 ?????

OpenStudy (aayushi.somani):

i think 2 is on the denominator of the whole and not of the power..! @RadEn

OpenStudy (whpalmer4):

I think more likely \[64^{\frac{3}{2}}\]given the size of \[\frac{64^3}{2}\]